Q1) Suppose GDP is GH¢ 8 trillion, taxes are GH¢ 1.5 trillion, private saving is GH¢ 0.5 trillion and public saving is GH¢ 0.2 trillion. Assuming this economy is closed. Calculate

a) consumption

b) Government purchases

c) National saving

d) Investment


Q2) The following are the exchange rate between Ghana cedi (GH¢) and the US Dollar ( USD) at times T and T+1.

Exchange rate at T is GH¢ 4.00/USD

Exchange rate at T+1 is GH¢ 5.00/USD.

a) Find the rate of appreciation of the US Dollar against the Ghana cedi from T to T+1.

b) The rate of depreciation of the Ghana cedi against the US Dollar from T to T+1.

c) Are the answers in (a) and (b) the same or equivalent? Explain.





1
Expert's answer
2021-06-03T13:44:15-0400

question1

closed economy: Y=C+I+G

a)     Consumption

Y=C+I+G

8=C+0.7+1.3

C=8-2

C=6trillion

b)     Government purchases

Public savings =(T-G)

0.2=1.5-G

G=1.5-0.2

=1.3trillion

c)     National saving

Private saving + public saving (0.2 +0.5) =0.7trilion.

d)     Investment

National saving=I=Private saving + public saving

Which is (0.2 +0.5) =0.7trilion.

 

 

 

 










Question2

Appreciate - when a currency becomes more valuable relative to another currency; a currency appreciates when you need more of another currency to buy a single unit of a currency.

Depreciate-when the value of a currency decreases relative to another currency; a currency depreciates when you need less of another currency to buy a single unit of a currency.

 

Exchange rate A​= unit of currency A /# of units of currency

a). exchange rate T=5.0UD/4.0USD=2.25USD

b) exchange rate T+1=4.0USD/5.0USD=0.8USD

c). the answer in a and b cannot be same as it is depreciation and appreciation.